Consider the reaction MnO2 + 4HCI → MnCl2 + Cl2 + 2H2O. If the both the


reactants have a mass of 100 grams, which of the two will be used up first? How


many grams of MnCl2 will be produced after the reaction? (MnO2 =


86.94 g/mol, HC1 = 36.46 g/mol, MnCl2 = 125.8 g/mol)

1
Expert's answer
2022-03-23T09:54:02-0400

Calculate "\nu:"

"\nu(MnO_2) =m(MnO_2)/M(MnO_2)=100g/(86.94 g/mol)=1.15 mol"

"\nu(HCl) =m(HCl)/M(HCl)=100g/(36.46 g/mol)=2.74 mol"


1 mol of "MnO_2" reacts with 4 mol of "HCl" , 1.15 mol of "MnO_2" reacts with 4.6 mol of "HCl"

"MnO_2" in excess

HCl will be used up first - it limiting reagent.


According to the equation we get 1/4 mol of "MnCl_2" from 1 mol of HCl, therefore:

"m(MnCl_2)=\nu(MnCl_2)\cdot M(MnCl_2)=1/4\cdot \nu(HCl)\cdot M(MnCl_2)="

"=0.25\cdot 2.74 mol\cdot 125.8 g/mol=86.17g"